10. A type of thermometer has temperature readings at the freezing point of water that are normally distributed with a mean of 0°C and a standard deviation of 1.00°C. Find the temperature corresponding to P23. the 23rd percentile. 0 0.9893 O 0.23° O-2.84 O-0.74
Answer #1

From the normal distribution table P ( x < -0.74 ) = 0.23
z = (x - μ) / σ
-0.74 = (x-0) / 1
x= -0.74

Temperature corresponding to P23, the 23rd percentile is -0.74C